Jess and Jillian went to the St. Anthony’s feast together. The admission fee is $5.00. At the feast there are 15 different types of rides to go on and to go on a ride you need to have tickets. The cost for one ticket is $2.00. Jess went on all 15 rides and Jillian went on all 15 rides and went on three rides twice. How much money did they pay altogether since they entered the Feast?
Answer by stanbon(75887)

Jess and Jillian went to the St. Anthony’s feast together. The admission fee is $5.00. At the feast there are 15 different types of rides to go on and to go on a ride you need to have tickets. The cost for one ticket is $2.00. Jess went on all 15 rides and Jillian went on all 15 rides and went on three rides twice. How much money did they pay altogether since they entered the Feast?
---------------
Jess expenses: 5 + 15*2 = $35
Jill expenses: 5 + 15*2 + 3*2 = 35+6 = $41
---------------------------------------
Total expenses: 35 + 41 = $76
=======================================
Cheers,
Stan H.